### v2

@@ Abstract
In the absence of comprehensive or random testing throughout the COVID-19 pandemic, we have developed a proxy method for synthetic random sampling to estimate community-level viral incidence, based on viral RNA testing of asymptomatic patients who present for elective procedures within a hospital system. The approach collects routine testing data on SARS-CoV-2 exposure among outpatients and performs statistical adjustments of sample representation using multilevel regression and poststratification (MRP), a procedure that adjusts for nonrepresentativeness of the sample and yields stable small group estimates. We have developed an open-source, user-friendly MRP interface for public implementation of the statistical workflow. We illustrate the MRP interface with an application to track community-level COVID-19 viral transmission in the state of Michigan.
Keywords: Bayesian workflow, bias correction, subgroup estimation, surveillance system
@@ 1 Introduction
In this paper, we present a user-friendly interface with an automatic implementation of multilevel regression and poststratification (MRP, Gelman and Little, 1997) and demonstrate the application to track community-level COVID-19 viral transmission. Previous work has developed the foundation of such a proxy metric for COVID-19 tracking data collection and statistical adjustment of demographic representation (Covello et al., 2021; Si et al., 2022). The approach collects electronic health records (EHRs) on routine viral testing of patients who present for elective procedures within a hospital system and performs MRP to estimate actual viral trends. The findings in a diverse urban-suburban-rural setting in Indiana show that this model predicts the clinical burden of SARS-CoV-2 earlier and more accurately than the currently accepted metrics. In contrast, the official testing data fail to inform the surge of clinical burdens. We extend the previous work by improving the MRP method and tracking community-level transmission—across geographic areas and demographic subgroups—to monitor the epidemic over time as an operational surveillance system. The interface uses the statistical programming language Stan (Stan Development Team, 2024) to conduct Bayesian computation and model estimation, specifically with the R package cmdstanr (Gabry et al., 2024). The interface is available at https://github.com/mrp-interface/shinymrp for local installation and at https://mrpinterface.shinyapps.io/shinymrp/ for web-based implementation.
As a prediction approach to modeling the outcome measures with individual-level and group-level predictors, MRP has become increasingly popular for subgroup estimation. Originally applied to estimate state-level public opinions from sociodemographic subgroups using sample surveys, MRP has two key components: (1) multilevel regression for small group estimation by setting up a predictive model with a large number of covariates and regularizing with Bayesian prior specifications, and (2) poststratification to adjust for selection bias and correct for imbalances in the sample composition from the target population. Flexible modeling of study outcomes can capture complex data structures conditional on poststratification cells, which are determined by the cross-tabulation of categorical auxiliary variables and calibrate the sample discrepancy with population control information. Besides applications in social sciences, especially in election forecasting (e.g., Wang et al., 2015; Lauderdale et al., 2020), MRP has also shown promise in public health research (e.g., Zhang et al., 2015; Downes et al., 2018; Downes and Carlin, 2020; Si et al., 2020). Si (2025) show that the key to the success of MRP in applications is the inclusion of highly predictive covariates, and Li and Si (2024) discuss estimation approaches when the population distribution of the poststratification variables is unknown.
@@
The interface enhances traditional MRP by introducing two novel methodological extensions: estimation varying across time and at granular geographic levels. Current applications of MRP predominantly rely on cross-sectional data collected at a single time point, either through probability sample surveys (e.g., Zhang et al., 2015; Si et al., 2020) or nonprobability samples (e.g., Wang et al., 2015; Si, 2025). In contrast, our interface extends MRP to accommodate time-varying data, enabling the tracking of trends. Moreover, due to sample size constraints, existing MRP estimates are typically at the national or state level. Our interface overcomes these limitations by supporting analyses at multiple geographic resolutions, including ZIP codes, counties, and states. The interface is capable of modeling both binary and continuous outcomes. As a demonstration, we apply the interface to time-varying, ZIP-code-level data to monitor demographic and county-level COVID-19 viral transmission trends in Michigan.
@@ 2.1 Data
Our COVID-19 tracking approach collects EHRs of prospective surgical (and other invasive procedure) patients who are asymptomatic and have tested for acute SARS-CoV-2 infection before performing the procedure (Covello et al., 2021; Si et al., 2022). Upon the reopening of elective medical and surgical procedures after the initial COVID-19 outbreak in early 2020, all preoperative patients were uniformly required—per the American Society of Anesthesiology guidelines—to undergo surgical risk evaluation and testing for acute SARS-CoV-2 infection before any such procedures. This policy was implemented nationwide across the U.S. All elective patients were presumed asymptomatic because any individual reporting symptoms or recent exposure to the virus would have their procedure either canceled or deferred. Using a standardized protocol, all preoperative patients underwent polymerase chain reaction (PCR) testing for viral RNA four days prior to their scheduled procedure, with tests administered by health system staff and samples analyzed using the same system. This PCR testing protocol was maintained consistently throughout the study period. Additionally, a subset of patients, for whom preoperative blood testing was clinically indicated based on age, health status, or surgery type, were also screened for the presence of immunoglobulin G (IgG) to the SARS-CoV-2 nucleocapsid protein (IgG N), beginning May 1, 2020.
@@
In collaboration with hospital database managers and in compliance with HIPAA privacy regulations, we collected EHR data including PCR test results, test dates, sex, age, race, and five-digit ZIP codes. The group represents broad age, racial, and socioeconomic diversity, with its only explicit correlation to disease status being the selection for elective surgical procedures and absence of symptoms or known exposure. We assume that, within any demographic and geographic stratum, the ratio of asymptomatic to symptomatic SARS-CoV-2 infections remains constant. Accordingly, the incidence of asymptomatic infections should proportionally reflect community-wide viral incidence and can serve as a proxy for true incidence trends, though this ratio may change with the emergence of new viral variants and the level of acquired immunity over time. To the extent that healthcare use or other factors affect the selection and ratio, we expect much of this variation to be addressed through our model adjustments. We discuss the potential violation of these assumptions in Section 4. MRP adjusts the demographic (sex, age, and race) and geographic (five-digit ZIP code) distributions to the target population. The target population is defined as U.S. residents dwelling in the catchment area of the collected ZIP codes. The interface links the input patient EHR data with ZIP codes to census tract measures in the American Community Survey (ACS), the largest household survey of the U.S. population (U. S. Bureau of the Census, 2025), and uses the ACS aggregated summaries of sociodemographic and socioeconomic characteristics as geographic predictors at the ZIP level.
@@
Previous work has treated PCR test sensitivity and specificity as unknown parameters, incorporating information from prior studies and accounting for estimation uncertainty in final MRP estimates (Gelman and Carpenter, 2020; Covello et al., 2021; Si et al., 2022). In the interface, users can specify different sensitivity and specificity values; for our demonstration, we presume 70% clinical sensitivity and 100% specificity, consistent with the previous setting during the same study period (March 2020–October 2022) (Bendavid et al., 2021; Gelman and Carpenter, 2020; Covello et al., 2021; Si et al., 2022).
@@
We track viral infections on a weekly basis. Below, we first introduce the conventional MRP framework for cross-sectional data, followed by extensions to accommodate time-varying tracking.
@@ 2.2 MRP for cross-sectional data
where ${\rm male}_{i}$ is an indicator for men, $\alpha_{\rm a}^{\rm age}$ is the age effect, with a value of $a[i]$ for subject $i$, on the log-odds function of the probability of having a positive response, $\alpha_{\rm r}^{\rm race}$ is the racial effect, and $\alpha_{\rm s}^{\rm ZIP}$ is the ZIP-code-level effect. In the Bayesian framework, we assign hierarchical priors to varying intercepts as default:
@@
Here $\mbox{normal}_{+}(0,2.5)$ represents a half-normal distribution with the mean 0 and standard deviation 2.52.52.5 restricted to positive values. As we have ZIP-code-level predictors $\vec{Z}^{\rm ZIP}_{s}$, we need to build another model in which $\alpha_{\rm s}^{\rm ZIP}$ is the outcome of a linear regression with ZIP-code-level predictors:
@@
The interface allows users to specify alternative priors, including structured priors for high-order interaction terms (Si et al., 2020). We use the default normal priors as examples and discuss extensions to spatial modeling in Section 4.
[…]
To generate overall population or subgroup estimates, we combine model predictions within the poststratification cells—in the contingency table of sex, age, race, and ZIP—weighted by the population cell frequencies $N_{j}$, which are derived from the linked ACS data in our application. Additionally, users may choose to upload custom poststratification data for specific target populations (e.g., a different country, rather than the U.S.). If we write the expected outcome in cell $j$ based on model (2.2.1) as $\hat{\theta}_{j}$ in cell $j$, the population average from MRP is then:
@@ 2.3 MRP for time-varying data
[…]
@@
As an example, we assign normal priors to the ZIP-code-level and time-varying effects. The interface leverages Stan’s modeling capabilities to allow alternative prior choices and can be extended with advanced modeling, such as spatial priors (Si et al., 2015) for ZIP-code-level effects or time series priors (e.g., first-order autoregressive) for temporal effects. Alternative outcome models (e.g., negative binomial) can be specified to accommodate overdispersion. In our COVID-19 application, we did not find substantial differences in the examined group estimates with various outcome model and prior specifications, so we presented the results based on a binomial model with normal priors. We elaborate further on model extensions in Section 4.
@@ 3 Bayesian workflow with MRP
The interface implements an end-to-end Bayesian MRP workflow of statistical analyses, from data preprocessing, descriptive summaries, model fitting, diagnostics, to presentation of results, following the principles of Gelman et al. (2020). For illustration, we apply this process to COVID-19 tracking in Michigan and validate the findings in comparison with other studies.
@@ 3.1 Data preprocessing
The interface accepts input as either individual patient test records or aggregated summaries at the poststratification cell level. In the initial step, the interface reads and displays the input data. The data cleaning and linking process is illustrated by the flowchart in Figure 3.1. This workflow automatically imputes missing predictor values using observed frequency distributions, converts categorical variables to factors, and standardizes continuous predictors at the ZIP code level. The MRP integrates three data sources: (1) PCR test results, (2) poststratification cell counts from the ACS, and (3) ZIP-code-level predictors linked from the ACS.
The interface identifies the relevant measures in PCR test records from hospitals for the modeling step. Specifically, the input data frame must include columns representing key demographic, geographic, and temporal measures: sex, race, age, five-digit ZIP code, PCR test result, and result date.
@@
Balancing operational feasibility, timeliness and accuracy, patient records are linked to the 2021 five-year ACS dataset by residential ZIP codes, using the R package tidycensus (Walker et al., 2025). This linkage serves two purposes: (1) defining the target population as people living in the ZIP codes’ catchment area and deriving population counts for poststratification cells, and (2) incorporating area-level predictors of viral infection to adjust for geographic variation. While the ACS reports geography at the levels of census tracts, counties, and states, ZIP codes are defined by the U.S. Postal Service (USPS). We use the ZIP code crosswalk table released by the U.S. Department of Housing and Urban Development and USPS to link ZIP codes to census tracts (U.S. Department of Housing and Urban Development, 2023) and calculate the ZIP-code-level measures by aggregating all available tract-level measures weighted by tract population counts. We select the county with the most-overlapping residential addresses for one ZIP code as the ZIP-linked county. The catchment area covered by the list of residence address ZIP codes provided by the Michigan Medicine patients can cover multiple states, beyond Michigan. We filter the data geographically by first removing ZIP codes with five or fewer records and then states that constitute less than 1% of the remaining data.
We construct poststratification cells by cross-classifying sex, race, age, and ZIP code and obtain the population counts for these cells from weighted ACS sample distributions in the relevant catchment area. These counts are assumed to remain constant throughout the study period (2020–2022).
@@
The geographic predictors include both individual-level variables (such as education, employment, and income) and tract-level variables (including urbanicity and the Area Deprivation Index (ADI, Kind and Buckingham, 2018). These are aggregated to ZIP codes as follows: (1) urbanicity: the percentage of covered census tracts classified as urban, weighted by tract population; (2) college: the percentage of residents with an Associate’s degree or higher; (3) poverty: the percentage of residents with incomes below the poverty level in the past year; (4) employment: the percentage of the civilian labor force that is employed; (5) income: the population-weighted average of tract-level median household incomes over the past 12 months; and (6) ADI: the population-weighted average of tract-level ADI values across covered census tracts.

### v3

@@ Abstract
We present a novel Bayesian workflow for multilevel regression and poststratification (MRP), introducing extensions to time-varying data and granular geography and publicly available open-source computation tools, facilitating broad research adoption and reproducibility. In the absence of comprehensive or random testing throughout the COVID-19 pandemic, we have developed a proxy method for synthetic random sampling to estimate community-level viral incidence, based on viral RNA testing of asymptomatic patients who present for elective procedures within a hospital system. The approach collects routine testing data on SARS-CoV-2 exposure among outpatients and performs statistical adjustments of sample representation using MRP, a procedure that adjusts for nonrepresentativeness of the sample and yields stable small group estimates. We illustrate the MRP interface with an application to track community-level COVID-19 viral transmission in the state of Michigan.
Keywords: Bayesian workflow, bias correction, subgroup estimation, surveillance system, reproducibility
@@ 1 Introduction
We present a novel Bayesian workflow for multilevel regression and poststratification (MRP, gelman:little:97), which is built on a user-friendly open-source interface, and demonstrate the application to track community-level COVID-19 viral transmission. Previous work has developed the foundation of such a proxy metric for COVID-19 tracking data collection and statistical adjustment of demographic representation (mrp-covid21; mrp-covid22). The approach collects electronic health records (EHRs) on routine viral testing of patients who present for elective procedures within a hospital system and performs MRP to estimate actual viral trends. The findings in a diverse urban-suburban-rural setting in Indiana show that this model predicts the clinical burden of SARS-CoV-2 earlier and more accurately than the currently accepted metrics. In contrast, the official testing data fail to inform the surge of clinical burdens.
@@
Our new development enhances traditional MRP by introducing two novel methodological extensions: estimation varying across time and at granular geographic levels. Current applications of MRP predominantly rely on cross-sectional data collected at a single time point, either through probability sample surveys (e.g., Zhang15-mrp; prior-si2018) or nonprobability samples (e.g., wang:gelman14; mrp-si20). In contrast, this paper extends MRP to accommodate time-varying data, enabling the tracking of trends. Moreover, due to sample size constraints, existing MRP estimates are typically at the national or state level. Our interface overcomes these limitations by supporting analyses at multiple geographic resolutions, including ZIP codes, counties, and states. The interface is capable of modeling both binary and continuous outcomes. As a demonstration, we apply the interface to time-varying, ZIP-code-level data to monitor demographic and county-level COVID-19 viral transmission trends in Michigan. We generate granular, timely and reproducible findings to direct resource allocation, guide surge prediction and inform rapid public health responses and evidence-based policy making.
As a prediction approach to modeling the outcome measures with individual-level and group-level predictors, MRP has become increasingly popular for subgroup estimation. Originally applied to estimate state-level public opinions from sociodemographic subgroups using sample surveys, MRP has two key components: (1) multilevel regression for small group estimation by setting up a predictive model with a large number of covariates and regularizing with Bayesian prior specifications, and (2) poststratification to adjust for selection bias and correct for imbalances in the sample composition from the target population. Flexible modeling of study outcomes can capture complex data structures conditional on poststratification cells, which are determined by the cross-tabulation of categorical auxiliary variables and calibrate the sample discrepancy with population control information. Besides applications in social sciences, especially in election forecasting (e.g., wang:gelman14; Yougov:mrp20), MRP has also shown promise in public health research (e.g., Zhang15-mrp; Carlin:AJE2018; carlin:AJE20; prior-si2018). mrp-si20 show that the key to the success of MRP in applications is the inclusion of highly predictive covariates, and emrp21 discuss estimation approaches when the population distribution of the poststratification variables is unknown.
The new workflow improves the MRP method with an application of tracking community-level transmission—across geographic areas and demographic subgroups—to monitor the epidemic over time using the developed interface as an operational surveillance system. Leveraging the statistical programming language Stan (stan-software:2021) to conduct Bayesian computation and model estimation, specifically with the R package cmdstanr (cmdstanr), the user-friendly interface promotes methodological transparency, reproducibility, and standardization for public health and social science research. The interface is accessible via both web-based and local platforms using the R package shinymrp (shinymrp-cran), offering flexible tools to facilitate analyses for diverse users. By integrating data preparation, model fitting, diagnostic checks, and graphical reporting into a comprehensive end-to-end pipeline, the interface enables researchers and practitioners to standardize MRP analyses across studies and datasets. This feature substantially lowers technical barriers, allowing reproducible application of advanced statistical methods and fostering comparability in population health research. We illustrate model use and interface capabilities with a real-world application tracking community-level COVID-19 viral transmission in Michigan. Health agencies can use subgroup-specific incidence estimates to allocate testing resources and direct targeted interventions to communities with greater vulnerability. Tracking epidemiological trends in real time through routine hospital data provides actionable guidance for surge prediction and rapid public health responses. While our primary application centers on community-level COVID-19 monitoring, the MRP workflow developed here is broadly applicable across a range of public health surveillance contexts.
@@ 2.1 Data
Our COVID-19 tracking approach collects electronic health records of prospective surgical (and other invasive procedure) patients who are asymptomatic and have tested for acute SARS-CoV-2 infection before performing the procedure (mrp-covid21; mrp-covid22). Upon the reopening of elective medical and surgical procedures after the initial COVID-19 outbreak in early 2020, all preoperative patients were uniformly required—per the American Society of Anesthesiology guidelines—to undergo surgical risk evaluation and testing for acute SARS-CoV-2 infection before any such procedures. This policy was implemented nationwide across the U.S. All elective patients were presumed asymptomatic because any individual reporting symptoms or recent exposure to the virus would have their procedure either canceled or deferred. Using a standardized protocol, all preoperative patients underwent polymerase chain reaction (PCR) testing for viral RNA four days prior to their scheduled procedure, with tests administered by health system staff and samples analyzed using the same system. This PCR testing protocol was maintained consistently throughout the study period. Additionally, a subset of patients, for whom preoperative blood testing was clinically indicated based on age, health status, or surgery type, were also screened for the presence of immunoglobulin G (IgG) to the SARS-CoV-2 nucleocapsid protein (IgG N), beginning May 1, 2020.
@@
In collaboration with hospital database managers and in compliance with HIPAA privacy regulations, we collected EHR data including PCR test results, test dates, sex, age, race, and five-digit ZIP codes. The group represents broad age, racial, and socioeconomic diversity, with its only explicit correlation to disease status being the selection for elective surgical procedures and absence of symptoms or known exposure. We assume that, within any demographic and geographic stratum, the ratio of asymptomatic to symptomatic SARS-CoV-2 infections remains constant. Accordingly, the incidence of asymptomatic infections should proportionally reflect community-wide viral incidence and can serve as a proxy for true incidence trends, though this ratio may change with the emergence of new viral variants and the level of acquired immunity over time. To the extent that healthcare use or other factors affect the selection and ratio, we expect much of this variation to be addressed through our model adjustments. We discuss the potential violation of these assumptions in Section 4. MRP adjusts the demographic (sex, age, and race) and geographic (five-digit ZIP code) distributions to the target population. The target population is defined as U.S. residents dwelling in the catchment area of the collected ZIP codes. We link the input patient EHR data with ZIP codes to census tract measures in the American Community Survey (ACS), the largest household survey of the U.S. population (acs2021), and use the ACS aggregated summaries of sociodemographic and socioeconomic characteristics as geographic predictors at the ZIP level.
@@
Previous work has treated PCR test sensitivity and specificity as unknown parameters, incorporating information from prior studies and accounting for estimation uncertainty in final MRP estimates (specificity:gelman20; mrp-covid21; mrp-covid22), which we recommend for use when their values are undetermined. The interface allows for specifying different sensitivity and specificity values; for our demonstration, we presume 70% clinical sensitivity and 100% specificity, consistent with the previous setting during the same study period (March 2020–October 2022) (bendavid2021covid; specificity:gelman20; mrp-covid21; mrp-covid22).
@@
We track viral infections on a weekly basis. Below, we first introduce the conventional MRP framework for cross-sectional data, followed by extensions to accommodate time-varying data features.
@@ 2.2 MRP for cross-sectional data
where ${\rm male}_{i}$ is an indicator for men, $\alpha_{\rm a}^{\rm age}$ is the age effect, with a value of $a[i]$ for subject $i$, on the log-odds function of the probability of having a positive response, $\alpha_{\rm r}^{\rm race}$ is the racial effect, and $\alpha_{\rm s}^{\rm ZIP}$ is the ZIP-code-level effect. In the Bayesian framework, we assign hierarchical priors to varying intercepts as a default setting:
@@
Here $\mbox{normal}_{+}(0,2.5)$ represents a half-normal distribution with the mean 0 and standard deviation $2.5$ restricted to positive values. As we have ZIP-code-level predictors $\vec{Z}^{\rm ZIP}_{s}$, we need to build another model in which $\alpha_{\rm s}^{\rm ZIP}$ is the outcome of a linear regression with ZIP-code-level predictors:
@@
Alternative prior distributions, including structured priors for high-order interaction terms (prior-si2018) and spatial prior accounting for geospatial correlation, can also be specified. We use the default setting in the interface with normal priors given in (2.2) and (2.2.3) as examples and discuss extensions to other settings in Section 4.
Because (2.2.1) assumes that the people in the same poststratification cell share the same response probability, we can replace the microdata with cellwise aggregates and employ a binomial model for the sum of the responses in cell $j$ as $y^{*}_{j}\sim\textrm{binomial}(n_{j},\theta_{j})$, where $n_{j}$ is the sample cell size and $\theta_{j}=\textrm{logit}^{-1}(\beta_{1}+\beta_{2}{\rm male}_{j}+\alpha_{\rm a[j]}^{\rm age}+\alpha_{\rm r[j]}^{\rm race}+\alpha_{\rm s[j]}^{\rm ZIP})$ using the cellwise effects of all factors. The input data for analysis can thus be either microdata or cellwise aggregates.
To generate overall population or subgroup estimates, we combine model predictions within the poststratification cells—in the contingency table of sex, age, race, and ZIP—weighted by the population cell frequencies $N_{j}$, which are derived from the linked ACS data in our application. Additionally, we may choose custom poststratification data for specific target populations (e.g., a different country, rather than the U.S.). If we write the expected outcome in cell $j$ based on model (2.2.1) as $\hat{\theta}_{j}$ in cell $j$, the population average from MRP is then:
@@ 2.3 MRP for time-varying data
We fit a binomial model for $y^{*}_{j}$, $y^{*}_{j}\sim\textrm{binomial}(n_{j},p_{j})$ with a logistic regression for $\pi_{j}$ with covariates—sex, age, race, ZIP codes, and time in weeks—to allow time-varying incidence in the multilevel model.
@@
As an example, we assign normal priors to the ZIP-code-level and time-varying effects. The interface leverages Stan’s modeling capabilities to allow alternative prior choices and can be extended with advanced modeling (BNFP:SI15), such as spatial priors for ZIP-code-level effects or time series priors (e.g., first-order autoregressive) for temporal effects. Alternative outcome models (e.g., negative binomial) can be specified to accommodate overdispersion. In our COVID-19 application, we did not find substantial differences in the examined group estimates with various outcome model and prior specifications, so we presented the results based on a binomial model with normal priors. We elaborate further on model extensions in Section 4.
@@ 3 Bayesian workflow with MRP
The interface implements an end-to-end Bayesian MRP workflow of statistical analyses, from data preprocessing, descriptive summaries, model fitting, diagnostics, to presentation of results, following the principles of workflow:gelman21. For illustration, we apply this process to COVID-19 tracking in Michigan and validate the findings in comparison with other studies.
@@ 3.1 Data preprocessing
The input data can be either individual patient test records or aggregated summaries at the poststratification cell level. The data cleaning and linking process is illustrated by the flowchart in Figure 3.1. This workflow automatically imputes missing predictor values using observed frequency distributions, converts categorical variables to factors, and standardizes continuous predictors at the ZIP code level. The MRP integrates three data sources: (1) PCR test results, (2) poststratification cell counts from the ACS, and (3) ZIP-code-level predictors linked from the ACS.
@@
The PCR test records from hospitals include key demographic, geographic, and temporal measures: sex, race, age, five-digit ZIP code, PCR test result, and result date, for the modeling. Balancing operational feasibility, timeliness and accuracy, patient records are linked to the 2021 five-year ACS dataset by residential ZIP codes, using the R package tidycensus (walker2021package). This linkage serves two purposes: (1) defining the target population as people living in the ZIP codes’ catchment area and deriving population counts for poststratification cells, and (2) incorporating area-level predictors of viral infection to adjust for geographic variation. While the ACS reports geography at the levels of census tracts, counties, and states, ZIP codes are defined by the U.S. Postal Service (USPS). We use the ZIP code crosswalk table released by the U.S. Department of Housing and Urban Development and USPS to link ZIP codes to census tracts (hud_usps_ziptract_crosswalk_2022) and calculate the ZIP-code-level measures by aggregating all available tract-level measures weighted by tract population counts. We select the county with the most-overlapping residential addresses for one ZIP code as the ZIP-linked county. The catchment area covered by the list of residence address ZIP codes provided by the Michigan Medicine patients can cover multiple states, beyond Michigan. We filter the data geographically by first removing ZIP codes with five or fewer records and then states that constitute less than 1% of the remaining data. We construct poststratification cells by cross-classifying sex, race, age, and ZIP code and obtain the population counts for these cells from weighted ACS sample distributions in the relevant catchment area. These counts are assumed to remain constant throughout the study period (2020–2022).
@@
The geographic predictors include both individual-level variables (such as education, employment, and income) and tract-level variables (including urbanicity and the Area Deprivation Index (ADI, kind2018making). These are aggregated to ZIP codes as follows: (1) urbanicity: the percentage of covered census tracts classified as urban, weighted by tract population; (2) college: the percentage of residents with an Associate’s degree or higher; (3) poverty: the percentage of residents with incomes below the poverty level in the past year; (4) employment: the percentage of the civilian labor force that is employed; (5) income: the population-weighted average of tract-level median household incomes over the past 12 months; and (6) ADI: the population-weighted average of tract-level ADI values across covered census tracts.